# Unexplainability of Artificial Intelligence Judgments in Kant’s Perspective

### v2

@@ Abstract
Kant’s Critique of Pure Reason, a major contribution to the history of epistemology, proposes a table of categories to elucidate the structure of the a priori principle of human judgment. The technology of artificial intelligence (AI), based on functionalism, claims to simulate or replicate human judgment. To assess this claim, it is necessary to study whether AI judgment possesses the characteristics of human judgment. This paper argues that AI judgments exhibit a form that cannot be understood in terms of the characteristics of human judgments according to Kant. Because the characteristics of judgment overlap, we can call this AI’s uncertainty. Then, I show that concepts without physical intuitions are not easy to explain when their functions are shown through vision. Finally, I illustrate that even if AI makes sentences through subject and predicate in natural language, which are components of judgment, it is difficult to determine whether AI understands the concepts to the level humans can accept. This shows that it is questionable whether the explanation through natural language is reliable.
@@ Introduction
Philosophy has explored the concept of human intelligence for thousands of years, and the study of intelligence is a traditional philosophical topic. Logic attempts to explain the structure of knowledge, and epistemology seeks to understand how knowledge is obtained. In computer science, the powerful learning capabilities of deep learning (LeCun, Bengio, and Hinton 2015) have attracted attention. In particular, artificial neural networks (ANNs) (LeCun, Bengio, and Hinton 2015) are commonly used to mimic human intelligence. However, the information that has been produced with respect to the type of judgment used by artificial intelligence (AI) remains insufficient. Indeed, the nature of the judgment made by AI is often considered a black box (Adadi and Berrada 2018). Philosophy, with its tools to explain and analyze intelligence, should investigate the mimicry of the intelligence of ANNs.
First, I examine this topic broadly using Kant’s four categories of judgment to show that unexplainability can occur in all of them. Then, explainable AI technology is introduced. I then discuss the SoftMax function (Sharma, Sharma, and Athaiya 2017) that can force AI judgments to make possibility judgments. Next, the limitations of vision-based explainable AI technology are examined based on the fact that concepts whose functions are revealed in vision but have no physical intuitions are difficult to explain. The self and self-consciousness are taken as examples. Finally, I show that even by using natural language processing technology that reveals subjects and predicates for judgments a machine makes, it is difficult to determine whether a machine understands the judgments it makes at a level that a human would accept. Therefore, this shows that explanation through natural language processing is not reliable.
@@
(Kant 1855), which translated Kant’s book into English, is being used in this work because Kant’s Critique of Pure Reason was not originally written in English. This paper investigates AI judgments. For this, it is necessary to understand what judgment is. Judgments are divided in philosophy into subjects and predicates. For example, in the sentence “Socrates is wise,” “Socrates” is the subject, and “is wise” is the predicate. Basic information about judgment can be obtained in Kant’s discussion of analytic and synthetic judgment (Kant 1855, p. 7).
In all judgments wherein the relation of a subject to the predicate is cogitated, (I mention affirmative judgments only here; the application to negative will be very easy,) this relation is possible in two different ways. Either the predicate B belongs to the subject A, as somewhat which is contained (though covertly) in the conception A; or the predicate B lies completely out of the conception A, although it stands in connection with it. In the first instance, I term the judgment analytical, in the second, synthetical. Analytical judgments (affirmative) are therefore those in which the connection of the predicate with the subject is cogitated through identity; those in which this connection is cogitated without identity, are called synthetical judgments. The former may be called explicative, the latter augmentative judgments; because the former add in the predicate nothing to the conception of the subject, but only analyse it into its constituent conceptions, which were thought already in the subject, although in a confused manner; the latter add to our conceptions of the subject a predicate which was not contained in it, and which no analysis could ever have discovered therein. For example, when I say, “All bodies are extended,” this is an analytical judgment. For I need not go beyond the conception of body in order to find extension connected with it, but merely analyse the conception, that is, become conscious of the manifold properties which I think in that conception, in order to discover this predicate in it: it is therefore an analytical judgment. On the other hand, when I say, “All bodies are heavy,” the predicate is something totally different from that which I think in the mere conception of a body. By the addition of such a predicate therefore, it becomes a synthetical judgment. Judgments of experience, as such, are always synthetical.
In his exploration of the possibility of synthetic a priori judgments, Kant intricately maps the capabilities of intelligence, and he illustrates and defines the forms of judgment (Kant 1855, p. 58).
If we abstract all the content of a judgment, and consider only the intellectual form thereof, we find that the function of thought in a judgment can be brought under four heads, of which each contains three momenta.
Based on this claim, he presents Table 1 describing the features of judgments. Only through these can our intellect understand and think.
Quantity Universal Particular Singular
@@
Modality Problematical Assertorical Apodeictical
Table 1: Logical function of the understanding in judgments (Kant 1855, p. 58).
Table 1 can make us assign each judgment to a specific category. For example, the proposition “All the values of 36 in pixels made by an infrared camera might indicate an actual human,” relates to universal, affirmative, categorical, and problematical.
@@
Artificial intelligence
A. M. Turing provided a striking test to assess the question of whether machines can think which is now known as the Turing test (French 2000; Turing 2009). According to this, if a machine provides responses that are indistinguishable from the way that a human would respond, then the two should not be differentiated. This aligns with functionalism (Levin 2004) in the philosophy of mind.
Computer scientists generally and unwittingly adopt functionalist perspectives on the human mind. Many of these scholars contend that when computers perform human-like functions, we can recognize them as AI.
As a prominent AI algorithm, ANNs incorporate interconnected artificial neurons. Generally, ANNs are trained to generate correct output values that correspond to certain given input values. The significance of ANNs as a representative algorithm is evident in studies demonstrating their ability to efficiently train approximate solutions for all functions (Csáji et al. 2001; Hornik, Stinchcombe, and White 1989; Schäfer and Zimmermann 2006; Zhou 2020). This implies that ANNs can find all functions, and therefore, it can be assumed that all functions can be computed by using ANNs.
I would like to emphasize that the process of forward propagation, through which artificial intelligence makes judgments, is completely mathematically explained. However, what judgments are made is not explained like human judgment. During their training, all of the error contributions generated by neurons are calculated using a backpropagation algorithm (LeCun, Bengio, and Hinton 2015), which enables the ANN to learn by minimizing errors. This backpropagation algorithm provides information in a readable form using numbers. However, this form cannot be considered to be the equivalent of a human judgment. We know why AI makes judgments in numerical format, but we do not know expressions resembling human judgment. The field that aims to interpret AI judgments like human judgments is called explainable AI (Adadi and Berrada 2018).
@@
When AI makes a judgment, it produces numerical outputs. For instance, in image classification, let us say that if an image contains a face, the AI outputs 1; otherwise, it outputs 0. However, the determination of whether the judgment corresponds to the form of “Universal,” “Particular,” or “Singular” in Table 1 is challenging. For instance, when an ANN learns to classify a human face, people may require additional information to distinguish among statements such as “This image has a human face,” “Some of these types of images have a human face,” and “All these types of images have a human face.” Human beings’ expressions of their judgments are made in their exact form with natural language. However, AI outputs are represented as numbers. Therefore, when AI makes a specific judgment, we cannot classify its form in quantity in Table 1.
In logic, only “Affirmative” (A is B) and “Negative” (A is not B) are deemed necessary. However, Kant distinguishes “Infinite” from “Negative” (A is (not B)) (Kant 1855, p. 59).
If I say of the soul, “It is not mortal,”—by this negative judgment I should at least ward off error. Now, by the proposition, “The soul is not mortal,” I have, in respect of the logical form, really affirmed, inasmuch as I thereby place the soul in the unlimited sphere of immortal beings. Now, because, of the whole sphere of possible existences, the mortal occupies one part, and the immortal the other, neither more nor less is affirmed by the proposition, than that the soul is one among the infinite multitude of things which remain over, when I take away the whole mortal part. But by this proceeding we accomplish only this much, that the infinite sphere of all possible existences is in so far limited, that the mortal is excluded from it, and the soul is placed in the remaining part of the extent of this sphere. But this part remains, notwithstanding this exception, infinite, and more and more parts may be taken away from the whole sphere, without in the slightest degree thereby augmenting or affirmatively determining our conception of the soul.
@@
It is not clear whether ANN outputs correspond to forms of “Affirmative,” “Negative,” or “Infinite.” For example, let us take again AI that is trained to output 1 when an image contains a human face and 0 when it does not. The judgments it makes during training with a value of 0 could be interpreted as “This image has a nonhuman face” or “This image does not have a human face.”
@@
Let me examine several cases in detail. While “This image has a human face,” “This image does not have a human face,” and “This image has a nonhuman face” are all plausible interpretations, some are ruled out as a result of the characteristics of the task. First, we can imagine that the AI only uses the “Affirmative” form of a judgment when it outputs 1 or 0. When it outputs 1, it states, “This image has a human face.” However, when it outputs 0, no predicate can be affirmed, except for the affirmation of the negative predicate itself. Hence, we conclude that ANN utilizes a combination of the “Affirmative” and “Infinite” forms of judgment. Second, we can imagine that the AI makes judgments that only take the “Negative” form. Here, an output of 0 would mean that “This image does not have a human face,” and an output of 1 would not produce any corresponding predicate, with the exception of the “Affirmative” form. Hence, judgments are only possible involving a combination of “Affirmative” and “Negative” judgments. The use of only the “Infinite” form does not provide a solution. Additionally, with the exception of the two combinations of judgments determined above, a combination of the “Infinite” and “Negative” judgments would also seem impossible. If the ANN utilizes a combination of the “Affirmative” and “Infinite” forms of judgment, the question then arises: Is this type of learning acceptable? The ANN could be affirming that something is a nonhuman face but lacks further specification.
From a logical point of view, the “Infinite” form is identical to the “Affirmative” form, and previous studies have found that ANNs perform their tasks effectively: They should output true values where we intended the output to be true and false where we intend the output to be false. However, due to the differences in content according to the quality form of the judgments, an ANN may decide not to make judgments in the way that we intend. For instance, it could learn to make judgments on an entirely different basis from humans. An ANN may process this in a mixed way, utilizing both “Negative” and “Infinite” forms.
@@
Considering Kant’s relation of judgment, this case presents even more difficult challenges than before. As summarized in Table 1, “Categorical” represents judgments on two concepts, “Hypothetical” indicates the relation between the two judgments, and “Disjunctive” signifies the relation between multiple judgments. For example, let’s consider a model that classifies objects as either a car, a dog, or a cat. It’s not always clear whether the model’s judgment is based on a single proposition, a causal relationship, or multiple propositions. The judgment could simply be “This image has a dog.” It could also be “This image has a dog because the image has a specific color.” Also, it could be “It is not a car and cat. I don’t know what a dog is. But it has to be a dog anyway. Therefore, there must be a dog.” In this last case, AI can make an accurate judgment without even possessing a full understanding of the concept of a dog. If this kind of learning occurs, it suggests that the learning process doesn’t align with our intention. Considering that AI is typically trained to classify objects within specific predefined classes, this phenomenon is quite common. Even when employing various technologies to elucidate decisions by highlighting the dog’s location in an image, it is uncertain whether AI comprehends the concept of a dog. This uncertainty arises because the AI might have made this determination based on distinguishing it from cats and cars.
[…]
$\vee¬(if ¬A, then ¬B)$
if (if A, then B), then ¬(if ¬A, then ¬B)
Table 2: Solution for the XOR problem.
Figure 1: Multilayer perceptrons solving the XOR problem.
Figure 1 shows that, although the value of true is defined as 1 and false as 0, there is no way to know whether AI has solved a logic problem or a classification problem. ANNs can solve the problem as a logical problem, as shown in Table 2, or as a problem of classification. Thus, we observe that unexplainability can exist in any of the four categories of judgment. Inspired by the uncertainty principle in quantum mechanics (Busch, Heinonen, and Lahti 2007), we can call this AI’s uncertainty. In this case, all forms of judgment overlap.
In the field of explainable AI research, various technologies are being studied (Adadi and Berrada 2018; Selvaraju et al. 2017). I believe that by developing technologies that provide insights into how AI judgments are made, we can achieve some level of explainability. Computer scientists have begun revealing which part of an image AI focuses on when making judgments. These tools assist us in inferring the subject and predicate. However, neither a subject nor a predicate is explicitly stated.
Among the four categories that are shown in Table 1, modality seems easier to put into explainable form than any of the others. This is because of the SoftMax function (Sharma, Sharma, and Athaiya 2017), which is applied to the final layer, ensuring that the cumulative sum of all options equals 1. This makes outputs satisfy the definition of probability. For example, when looking at an image and determining whether there is a dog or a cat, the SoftMax function makes it possible to indicate a 63% probability that there is a dog and a 37% probability that there is a cat. According to Kant, the modality form has the following characteristics (Kant 1855, pp. 60–61).
@@
The modality of judgments is a quite peculiar function, with this distinguishing characteristic, that it contributes nothing to the content of a judgment (for besides quantity, quality, and relation, there is nothing more that constitutes the content of a judgment), but concerns itself only with the value of linking verb in relation to thought in general.

### v4

@@ Abstract
Kant’s Critique of Pure Reason, a major contribution to the history of epistemology, proposes a table of categories to elucidate the structure of the a priori principles underlying human judgment. Artificial intelligence (AI) technology, grounded in functionalism, claims to simulate or replicate human judgment. To evaluate this claim, it is necessary to examine whether AI judgments exhibit the essential characteristics of human judgment. This paper investigates the unexplainability of AI judgments through the lens of Kant’s theory of judgment. Drawing on Kant’s four logical forms—quantity, quality, relation, and modality—this study identifies what may be called AI’s uncertainty, a condition in which different forms of judgment become entangled. In particular, with regard to modality, this study argues that the SoftMax function forcibly reframes AI judgments as possibility judgments. Furthermore, since complete definitions in natural language are impossible, words are, by their very nature, ultimately unexplainable; therefore, a fully complete functional implementation is theoretically unattainable.
@@ Introduction
Philosophy has explored the concept of human intelligence for thousands of years, and the study of intelligence is a traditional philosophical topic. For example, logic attempts to explain the structure of knowledge, and epistemology seeks to understand how knowledge is obtained. In computer science, the powerful learning capabilities of deep learning (LeCun, Bengio, and Hinton 2015) have attracted attention. In particular, artificial neural networks (ANNs) (LeCun, Bengio, and Hinton 2015) are commonly used to mimic human intelligence. These are models that simply mimic the functions of neurons in the brain, making them mathematically tractable. From a mathematical perspective, these models are designed to perform desired intelligent functions by processing information through matrix multiplication involving many matrices. However, information regarding the type of judgment employed by AI remains insufficient. Indeed, the nature of the judgment made by AI is often considered a black box (Adadi and Berrada 2018). In the AI community, the term “black box” indicates that the reason behind ANNs’ judgment is not presented in a manner comprehensible to humans. This phenomenon has been referred to as unexplainability in the AI community, which has led to the emergence of explainable AI as a field of research. Going further, it seems necessary to analyze this unexplainability—revealed in an indeterminate manner—more precisely using conceptual tools from philosophy, which has long engaged with the nature of intelligence.
Is AI truly capable of making judgments?
It may be argued that AI does not actually make judgments. It is not human. Nevertheless, the issue cannot be easily dismissed. Even if the notion of judgment cannot be applied to AI, we are still inevitably compelled to interpret its outputs through the lens of human reasoning. For instance, consider an AI model that outputs 1 when it detects a dog in an image. This numerical output is not understood as “1” itself; rather, humans interpret it as statements such as “a dog is present in the image,” “there is a high probability that the image contains a dog,” or “a dog may be in the bottom left of the image.” This reveals that we impose meaning from our own perspective. Such interpretations are independent of whether the machine is truly engaging in the act of judgment.
However, if machines do make judgments, a more serious problem arises: this leads to the question of whether explainable AI technology has fundamental limitations. We have striven to understand AI judgments within the confines of our human intellect. Can AI judgments possess additional characteristics beyond those of human judgments? Even if such judgments can exist, it appears that we lack the means to fully comprehend them. In that case, explainable AI technology is reduced to a tool that merely approximates machine judgments which fundamentally lie outside the bounds of human cognitive structures. Regardless of technological progress, true explainability may remain forever out of reach. If human cognition has inherent limitations and machines are able to surpass them, AI may come to occupy a unique position as an entirely new kind of epistemic agent.
I think that Turing also pursued his research (Turing 1950) with a similar sense of problem awareness. He posed the question, “Can machines think?” but seemed to realize that it would be impossible to offer a universally valid and absolute proof to answer it. Thus, he shifted the direction of the question: even if the physical substance differs, what does it mean if, in terms of mental and functional capabilities, a machine and a human being are indistinguishable? He proceeded to develop his ideas from this new angle. It seems that Turing shifts his perspective from an ontological level to a functional one. Ontology is a branch of philosophy concerned with what exists in the world—it deals with the nature of being itself. For example, one might ask, “Does the number 1 exist?” and “Does color actually exist, or is it dependent on human perception?”
One approach that seeks to address the nature of the mind at a structural level is functionalism (Levin 2004):
However, functionalism faces serious objections, particularly regarding whether functionality alone can account for genuine understanding. For instance, Searle’s Chinese Room argument (Searle 1980) demonstrates that the mind cannot be fully reduced to mere functionality. Simply put, just as one cannot be said to understand Chinese merely by receiving Chinese input and producing meaningful task outputs based on rules without knowing any Chinese, the mind cannot be reduced to a mere collection of functions. However, this is a study conducted at the ontological level, aiming to understand what the mind itself is. In other words, it is an analysis of the nature of the human mind. Given these ontological challenges, it becomes clear why AI researchers tend to adopt a purely functional perspective. From the perspective of implementing the functions of the mind, however, we do not necessarily need to know what the mind itself is. AI researchers are focusing on implementing the functions of the mind.
Observing the behavior of AI researchers today, they strive to artificially create entities that can be called “intelligence” and compete over how closely their creations resemble actual intelligence. When explaining AI systems that recognize and classify dogs, researchers often use terms related to intelligence such as judgment, pattern recognition, segmentation, and classification. However, when ordinary people look at such AI and say, “There is no difference from humans,” “I see self-consciousness in it,” researchers strongly refute such claims by insisting that AI is merely the result of mathematical computations. This contradiction is similar to the following: claiming “Truth is relative to social, cultural, and historical contexts,” yet demanding that this very claim be accepted as a universally valid truth, independent of any context.
Thus, this contradiction must be addressed first. AI researchers primarily focus on how to implement functionalities that resemble human intelligence. Therefore, even if their creations only approximate intelligence to a very small degree, they still call them AI. Every year, hundreds of papers are published in numerous journals to report new findings on neural networks in biology. Although we do not have a complete understanding of neural networks, we have made the definition of neural networks from our limited knowledge. Can such a definition be complete? Nevertheless, most people admit that ANNs have been created that resemble neural networks. In this sense, researchers are not discovering properties in an object, but rather assigning properties to it. In contrast, the general public observes the phenomena exhibited by AI and interprets them as the discovery of the properties of “intelligence” within it. For example, when seeing an AI that classifies images of dogs, people might say: “this system has properties of human intelligence.”
This broader context also forces researchers and me to reconsider the nature of the question being asked. In the same way, I, too, find myself inevitably having to shift my own question from an ontological framework to a functional one. Thus, the problem with my previous question arose because I tried to conceptually grasp and discover the essence of phenomena, rather than focusing on the functional aspect. We cannot immediately conclude that AI judgment is true judgment itself in the ontological sense. However, if the functionality of judgment has been implemented, it becomes necessary to philosophically examine how closely this functional implementation resembles actual human judgment.
Because AI judgment does not need to be ontologically identical to actual judgment, there is the possibility that it could possess new properties. Functionally, AI judgment is not a complete replica of human judgment. However, partial similarity is sufficient for it to be recognized as an AI judgment. And, from this partial resemblance, new properties may arise. Thus, a distinct space for the philosophy of AI emerges, precisely because a fundamental difference between actual intelligence and AI is inherently embedded.
Exploring the limitations of AI—identifying which functions can never fully replicate human capabilities—has become a vast topic within the philosophy of AI. At the same time, investigating whether certain functions might transcend or even overcome the limitations of human intelligence also constitutes a critical subject within this emerging philosophical domain. It is not necessary to contemplate abstract concepts such as ultimate knowledge. Intelligence encompasses a variety of functions, including multiplication—an area in which AI has already demonstrated superiority over humans. Also, it is possible for a system to perform a function without fully understanding or explaining the reasons behind its success. Thus, a degree of unexplainability is inherently embedded within the very execution of certain intelligent functions.
@@
Most readers of this paper are likely to be philosophers and AI researchers. AI is primarily studied within the fields of computer science and electrical engineering. Most people may not have studied Kant in depth. Conversely, philosophers typically do not receive an education in AI. Therefore, this paper has an obligation to provide sufficient background information to bridge these gaps. This study mainly explores the nature of judgment, and the two figures who made the most significant contributions to this topic in the history of philosophy are Aristotle and Kant. To draw an analogy, their roles in the philosophy of judgment are comparable to Newton and Einstein in the history of physics. While this study is based on Kant’s framework, it should be noted that Kant’s system is not considered perfect, and ongoing research continues in the field of philosophy. One might wonder why a study on the nature of judgment would draw upon Kant’s framework, which was formulated centuries ago. Although both figures passed away long ago, subsequent research has largely focused on overcoming the frameworks established by these two giants. This situation is analogous to how, although the Transformer algorithm (Vaswani et al. 2017) was introduced nearly a decade ago, most mainstream follow-up research still focuses merely on modifying that original algorithm. It is worth noting that the Transformer architecture still fundamentally relies on the linear layer (LeCun, Bengio, and Hinton 2015), a technique developed half a century ago. In this light, citing earlier philosophical frameworks such as Kant should not be hastily dismissed as outdated.
@@
The English translation (Kant 2024) is used in this work because Kant’s Critique of Pure Reason was not originally written in English. This paper investigates AI judgments, and for this purpose, it is necessary to clarify what a judgment is. Kant intricately maps and illustrates the features of judgment (Kant 2024, p. 34). He presents Table 1, which outlines the logical functions of judgment. Only through these forms can we engage in understanding and thought.
Table 1: Logical function of the understanding in judgments (Kant 2024, p. 34).
Quantity General Special Individual
@@
Modality Problematic Assertoric Apodictic
Each judgment corresponds to four categories—one from each domain in Kant’s Table. In summary, every judgment necessarily falls into one of three categories in terms of quantity: whether the subject refers to a single entity, multiple entities, or all entities. From the perspective of quality, a proposition must either affirm a predicate, deny it, or negate a negation. In terms of relation, judgments express the relation between two concepts, the relation between two propositions, or the relation among multiple propositions. In terms of modality, there are judgments of possibility, actuality, and necessity. Table 2 provides illustrative examples.
Table 2: Kant’s Types of Judgments with Examples
Type Example
General All humans are mortal.
Special Some animals are mammals.
Individual Socrates is wise.
Affirmative The apple is red.
Negative The apple is not red.
Infinite The apple is non-red (i.e., it belongs to the infinite set of things that are not red).
Categorical The ball is round.
Hypothetical If it rains, the ground gets wet.
Disjunctive It is either day or night.
Problematic It might be raining.
Assertoric It is raining.
Apodictic It must be raining, given the conditions.
@@
From Table 1, we can derive pure concepts of the understanding, which do not involve any empirical elements, and according to Kant, it is through these pure concepts that judgment is carried out. Table 3 shows it.
Table 3: Kant’s Table of Categories (Kant 2024, p. 37)
Category Group Categories
Quantity Unity, Multiplicity, Allness
Quality Reality, Negation, Limitation
Relation Substance and accident
Cause and effect
Community
Modality Possibility — impossibility
Existence — nonexistence
Necessity — contingency
All the concepts presented in Table 3 possess no empirical content whatsoever. For example, consider the concept of causality. Is causality something red? Is it large in size? We judge that one event is causally connected to another when we observe them repeatedly appearing in temporal succession. But, within the events themselves, no empirical material corresponding to “cause” can be found. Rather, it is only through such pure concepts that we are able to make causal judgments. Thus, these concepts belong to the form of thought required for experience.
For scientists, dealing with objects that lack empirical content might seem somewhat unfamiliar. However, in philosophy, there are fields that operate independently of experience, notably logic and metaphysics. Since no scientist denies the validity of logic, I will base my explanation on logic. If we open any standard textbook on logic, we find that propositions are typically represented by symbols such as A and B, and truth is inferred through logical symbols. What this means is that logic considers propositions only in terms of the aspects of thought that are entirely independent of content.
[…]
AI
ANNs are trained to generate correct output values that correspond to certain given input values. The significance of ANNs as a representative algorithm is evident in studies demonstrating their ability to efficiently train approximate solutions for all functions (Csáji et al. 2001; Hornik, Stinchcombe, and White 1989; Schäfer and Zimmermann 2006; Zhou 2020). This implies that ANNs can find all functions, and therefore, it can be assumed that all functions can be computed by using ANNs.
ANNs produce an output through matrix multiplication. This process is called a forward pass. Then, the difference between the output and the desired result is calculated, and in order to reduce this difference, the contribution of each matrix to the error during the forward pass is computed and propagated backward for updates. For example, if there are matrices A, B, and C, and the forward pass proceeds by sequentially multiplying an input by A, then by B, and then by C, C will be updated first according to the extent it contributed to the error, followed by B, and then A. The terms “forward” and “backward” were coined to describe this process.
I would like to emphasize that the process of forward propagation, through which AI makes judgments, is completely mathematically explained. However, what judgments are made is not explained like human judgment. During their training, all of the error contributions generated by matrices are calculated using a backpropagation algorithm (LeCun, Bengio, and Hinton 2015), which enables ANNs to learn by minimizing errors. This backpropagation algorithm provides information in a readable form using numbers. However, this form cannot be considered to be the equivalent of a human judgment. We know why AI makes judgments in numerical format, but we do not know expressions resembling human judgment.
@@
AI researchers naturally wonder why an AI system made a particular judgment. In doing so, they inquire not only into the surface outputs but into the underlying nature of the judgment itself. Once a system implements the functional structure of judgment, it is reasonable to assume that it also possesses, even if imperfectly, certain properties of judgment. Without such properties, the system’s activity would cease to qualify as judgment in any meaningful sense. Thus, once we attribute the concept of “judgment” even partially to AI behavior, the system must, to varying degrees, embody the properties associated with judgment. As previously discussed, the functional realization of judgment may vary in completeness, ranging from weak to fully developed forms.
Quantity, quality, and relation
@@
When AI makes a judgment, it produces numerical outputs. For instance, in image classification, let us say that if an image contains a face, the AI outputs 1; otherwise, it outputs 0. However, the determination of whether the judgment corresponds to the form of “Universal,” “Particular,” or “Singular” in Table 1 is challenging. For instance, when an ANN learns to classify a human face, people may require additional information to distinguish among statements such as “This image has a human face,” “Some of these types of images have a human face,” and “All these types of images have a human face.” Human beings’ expressions of their judgments are made in their exact form with natural language. However, AI outputs are represented as numbers. Therefore, when AI makes a specific judgment, we cannot classify its form in quantity in Table 1. Furthermore, it is unclear which pure concepts of the understanding listed in Table 3 should mediate the interpretation of AI outputs. In the following discussions, I will focus solely on the logical aspects without repeatedly mentioning the difficulty of mapping AI outputs to the pure concepts presented in Table 3.
@@
In logic, only “Affirmative” (A is B) and “Negative” (A is not B) are deemed necessary. However, Kant distinguishes “Infinite” from “Negative” (A is (not B)) (Kant 2024, pp. 34-35). It is not clear whether ANN outputs correspond to forms of “Affirmative,” “Negative,” or “Infinite.” For example, let us take again AI that is trained to output 1 when an image contains a human face and 0 when it does not. The judgments it makes during training with a value of 0 could be interpreted as “This image has a nonhuman face” or “This image does not have a human face.”
@@
Let me examine several cases in detail. While “This image has a human face,” “This image does not have a human face,” and “This image has a nonhuman face” are all plausible interpretations, some are ruled out as a result of the characteristics of the task. First, we can imagine that the AI only uses the “Affirmative” form of a judgment when it outputs 1 or 0. When it outputs 1, it states, “This image has a human face.” However, when it outputs 0, no predicate can be affirmed, except for the affirmation of the negative predicate itself. Hence, we conclude that ANN utilizes a combination of the “Affirmative” and “Infinite” forms of judgment. Second, we can imagine that the AI makes judgments that only take the “Negative” form. Here, an output of 0 would mean that “This image does not have a human face,” and an output of 1 would not produce any corresponding predicate, with the exception of the “Affirmative” form. Hence, judgments are only possible involving a combination of “Affirmative” and “Negative” judgments. The use of only the “Infinite” form does not provide a solution. Additionally, with the exception of the two combinations of judgments determined above, a combination of the “Infinite” and “Negative” judgments would also seem impossible. If the ANN utilizes a combination of the “Affirmative” and “Infinite” forms of judgment, the question then arises: Is this type of learning acceptable? The ANN could be affirming that something is a nonhuman face but lacks further specification. From a logical point of view, the “Infinite” form is identical to the “Affirmative” form, and previous studies have found that ANNs perform their tasks effectively: They should output true values where we intend the output to be true and false where we intend the output to be false. However, due to the differences in content according to the quality of the judgments, ANNs may decide not to make judgments in the way that we intend.
@@
Considering Kant’s relation of judgment, this case presents even more difficult challenges than before. As summarized in Table 1, “Categorical” represents judgments on two concepts, “Hypothetical” indicates the relation between the two judgments, and “Disjunctive” signifies the relation between multiple judgments. For example, let’s consider a model that classifies objects as either a car, a dog, or a cat. It’s not always clear whether the model’s judgment is based on a single proposition, two propositions, or multiple propositions. The judgment could simply be “This image has a dog.” It could also be “If this image exhibits certain learned features, then it is judged to contain a dog. In this case, the image satisfies those learned features. Therefore, the system judges that a dog is present.” Also, it could be “It is not a car and cat. I don’t know what a dog is. But it has to be a dog anyway. Therefore, there must be a dog.” A hypothetical judgment is first established by linking certain learned features to the classification of a dog. Upon confirming that the conditions are satisfied, the system finalizes the judgment in accordance with the hypothetical structure. In this last case, AI can make an accurate judgment without even possessing a full understanding of the concept of a dog. If this kind of learning occurs, it suggests that the learning process doesn’t align with our intention. Considering that AI is typically trained to classify objects within specific predefined classes, this phenomenon is quite common. Even when employing various technologies to elucidate decisions by highlighting the dog’s location in an image, it is uncertain whether AI comprehends the concept of a dog. This uncertainty arises because the AI might have made this determination based on distinguishing it from cats and cars.
Modality
According to Kant, the modality form has the following characteristics (Kant 2024, p. 35).
@@
The modality of judgments is a very special function of them, which has the distinguishing feature that it contributes nothing to the content of the judgment, (for apart from magnitude, quality, and proportion, there is nothing more that constituted the content of a judgment,) but only concerns the value of the copula in relation to thinking in general.
An example of this is illustrated in Table 4. Although the boiling point of water can vary depending on pressure, this factor is ignored here for simplicity. Moreover, whether the given proposition is actually true or false is not important in the present context.
Table 4: Expression of the same content and a single event in different modalities
Category Expression
(Fixed content with different modalities)
Problematic Water can boil at 100 degrees Celsius.
Assertoric Water is boiling at 100 degrees Celsius.
Apodictic Water must boil at 100 degrees Celsius.
(Single event with different modalities)
Problematic The object can break upon impact.
Assertoric The object is currently falling due to gravity.
Apodictic Upon analyzing the various intuitions given to me, it necessarily follows that an object falls due to gravity.
Kant’s theoretical framework emphasizes that while the content of a judgment remains fixed, the mode of its expression can vary. However, the focus of our discussion will be on the fact that a single phenomenon can be expressed through all three modalities. For example, consider a case where a computer solves an XOR logical problem. In human judgment, this is treated purely as a logical judgment, which is based on the modality of necessity. However, an AI system, which receives the input and output values as truth values, can approach the same problem either as a logical problem or as a classification task. It learns to produce the correct output depending on the input it receives. Thus, the outcome produced by AI can be: a logical consequence grounded in truth values (necessity), a judgment based on possibility by predicting probabilities through classification, or a classification judgment based on actuality by assigning the input to a specific class as a matter of fact (e.g., “It is A.”).
From this point, the probabilistic interpretation based on SoftMax function (Sharma, Sharma, and Athaiya 2017) begins to collapse. Softmax is a function that normalizes the outputs so that their sum becomes 1. Most AI researchers apply this function at the final stage, primarily for the purpose of probabilistic evaluation. For example, in a simple image classification task distinguishing between dogs and cats, applying Softmax produces an output like 37% for “dog” and 63% for “cat.” Although this is a common example in computer vision, Softmax is almost always used in natural language processing (NLP) as well. Even though the core algorithms for NLP tasks constantly evolve, the fundamental approach remains the same: the problem is reduced to predicting the next word based on the previous context. For instance, given the phrase “This morning I ate,” the model predicts “breakfast” as the most probable next word. This prediction is expressed as probabilities over all possible outputs, which can number in the tens of thousands depending on the model. Of course, we can design many functions other than Softmax that normalize outputs to sum to one. However, Softmax remains the most widely used. What I want to point out is that this is not a story limited to SoftMax. We often treat it as a kind of possibility judgment simply based on the condition that the sum of all possible cases equals 1. However, this reflects only a partial aspect of the full nature of possibility judgments, and it shows that we are not truly representing their complete characteristics. A possibility judgment does not necessarily have to be expressed in the form of a probability. For example, the statement “We can draw a circle” is a possibility judgment that is not quantified. In other words, we regard AI’s possibility judgments as such based on partial functional similarity.
I would like to briefly step away from the main topic to explain why NLP is handled in that way. The idea that meaning is determined by the way language is utilized is well expressed in Wittgenstein’s philosophy of language (Travis 1989) and plays an important role in the philosophy of language as well. It is worth noting that natural language processing is also designed with the view that the meaning of a word is determined by the existence of surrounding words (Devlin et al. 2018; Le and Mikolov 2014; Mikolov et al. 2013; Vaswani et al. 2017). It seems more convincing to say that they processed it in that way because the given data consisted of texts, rather than assuming they understood and implemented such a perspective from the philosophy of language.